OpenStudy (anonymous):

A Church spire in a form of a regular hexagonal pyramid whose base edge is 8ft and whose altitude is 75ft is to be painted at a cost of 22 dollars per sq. yard. What is the total cost?

OpenStudy (compassionate):

Is there any more information to go with the question?

OpenStudy (anonymous):

the altitude is 75ft

OpenStudy (compassionate):

Do you know the formula for a regular hexagonal pyramid

OpenStudy (compassionate):

Basically, find the area then multiply it by 22

OpenStudy (anonymous):

I kinda forgot

OpenStudy (compassionate):

surface area = 3ab + 3bs

OpenStudy (anonymous):

1/2 bh?

OpenStudy (compassionate):

That's for a triangle. I posted the formula bove

OpenStudy (anonymous):

Never seen that formula before

OpenStudy (anonymous):

Show me how it is done pls.

OpenStudy (compassionate):

I don't have time to finish this. @wio help him

OpenStudy (compassionate):

@KingGeorge

OpenStudy (kinggeorge):

Well from my brief research it seems like the surface area of a hexagonal pyramid is given by\[\frac{3\sqrt3}{2}a^2+3a\sqrt{h^2+\frac{3a^2}{4}}\] where \(a\) is the length of the base edge and \(h\) is the height.

OpenStudy (kinggeorge):

Does that formula look familiar at all?

OpenStudy (anonymous):

Very